# ARS: Automatic Routing Solver with Large Language Models

### v1

@@ Abstract
Real-world Vehicle Routing Problems (VRPs) are characterized by a variety of practical constraints, making manual solver design both knowledge-intensive and time-consuming. Although there is increasing interest in automating the design of routing algorithms, existing research has explored only a limited array of VRP variants and fails to adequately address the complex and prevalent constraints encountered in real-world situations. To fill this gap, this paper introduces RoutBench  [^1] https://github.com/Ahalikai/ARS-Routbench, a benchmark of 1,000 VRP variants derived from 24 attributes, for evaluating the effectiveness of automatic routing solvers in addressing complex constraints. Along with RoutBench, we present the Automatic Routing Solver (ARS), which employs Large Language Model (LLM) agents to enhance a backbone algorithm framework by automatically generating constraint-aware heuristic code, based on problem descriptions and several representative constraints selected from a database. Our experiments show that ARS outperforms state-of-the-art LLM-based methods and commonly used solvers, automatically solving 91.67% of common VRPs and achieving at least a 30% improvement across all benchmarks.
Vehicle Routing Problem, Large Language Model, Automatic Heuristic Design, Benchmark.
@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a fundamental Combinatorial Optimization Problem (COP) that plays a critical role in logistics, transportation, manufacturing, retail distribution, and delivery planning (Toth & Vigo, 2014; Liu et al., 2023). In these scenarios, the objective of VRPs is to efficiently allocate and plan vehicle routes to meet various requirements while minimizing overall routing costs. These requirements often include constraints such as vehicle capacities, time windows, and duration limits, resulting in numerous variants of VRPs in practical applications (Braekers et al., 2016). Existing heuristics are usually problem-specific. When the problem changes slightly (e.g., a minor modification to the requirements), a lot of effort is required for experts to redesign the heuristic to make it effective in solving the new problems (Vidal et al., 2013; Rabbouch et al., 2021; Errami et al., 2023).
Large Language Models (LLMs) have shown powerful reasoning and code-generation capabilities (Chen et al., 2021; Austin et al., 2021; Li et al., 2023). By integrating these functionalities, users can express their specific requirements in natural language, enabling the models to automatically design algorithms to address VRPs (Liu et al., 2024b). Most existing works primarily leverage the LLM to solve a small number of standard VRPs (Jiang et al., 2024; Huang et al., 2024) and cannot be applied to complex VRPs. Moreover, current methods rely on general solvers to optimize routes (Xiao et al., 2023; Zhang et al., 2024). However, solvers often struggle with standard modeling and exhibit low solving efficiency, making it difficult to handle complex constraints that arise in practical applications.
@@
  • We comprehensively validate our approach on commonly-used problems and RoutBench. The results show that ARS can automatically handle 91.67% of commonly-used problems, achieving at least a 30% improvement across all benchmarks compared to other LLM-based methods.
@@ 2 Problem Formulation
where $\mathbf{x}=\{x_{ij}\mid i,j\in\mathcal{V},\ i\neq j\}$ represents the set of decision variables, $x_{ij}$ is a binary variable that indicates whether the route from $i$ to $j$ is used. The feasible solution space $C$ is defined by constraints. In this paper, we consider VRPs with a variety of real-world constraints such as vehicle capacity, travel distance, and time windows, thereby extending the basic VRP, as seen in the Capacitated VRP (CVRP) (Toth & Vigo, 2014) and the VRP with Time Windows (VRPTW) (Solomon, 1987). Moreover, new constraints often emerge in real-world scenarios. For example, VRP variants related to vehicle capacity include Heterogeneous VRP (HVRP), which considers vehicles with different capacities (Lai et al., 2016), Multi-Product VRP (MPVRP), which addresses the need to transport multiple types of products (Yuceer, 1997), and dynamic demands (Powell, 1986). VRP variants that incorporate these real-world constraints are more prevalent and practically significant in real-world applications
@@ 3.1 Database
  • $C_{k}$: The constraint feasibility checking program. It checks whether a solution belongs to the feasible solution space described by $I_{k}$. The program is given a solution and returns ’True’ if the corresponding constraint is satisfied.
@@ 3.2 Constraint-Aware Heuristic
We first select relative constraints from the database. Then we sequenticaly generate the constraint checking and scoring programs for the target VRP variants given the selected constraints and the problem description. Finally, the constraint-aware heuristic is generated based on the designed constraint checking and scoring programs.
@@ 3.2.1 Constraint Selection
Given an input problem description $I$, we instruct LLM agents to automatically select a subset of relevant constraints $S$ from the database $D$ to provide reference information for subsequent constraint program generation. There can be two cases. In the first case, LLM agents select one or more relative constraint examples. In the second case, if no constraints are recognized as related to the input $I$, the base case $(I_{0},C_{0})$, with no additional contrasts, is selected.
This step is similar to Retrieval-Augmented Generation (RAG) methods that retrieve information from external sources (Zhang et al., 2023; Zhou et al., 2022), while our candidate reference information is a pre-defined database with representative constraints.
@@ 3.2.3 Constraint Scoring Program Generation
In practice, heuristic algorithms operate within a variable space that is typically much larger than the feasible solution space. This discrepancy presents significant challenges in identifying feasible solutions. To address this, heuristic methods may permit the presence of infeasible solutions during the search process, as this allows for the evaluation of potential improvements in solution quality (Deb, 2000; Máximo & Nascimento, 2021).
@@
3.2.4 Constraint-Aware Heuristic
We present the Constraint-Aware Heuristic (CAH) based on the constraint checking (Checker) and scoring programs (Scorer). As seen in Algorithm 1, the constraint handling method evaluates whether a new solution ${s}_{new}$ improves upon an old solution ${s}_{old}$. It first verifies whether ${s}_{new}$ is feasible (line 1). If ${s}_{new}$ is feasible and ${s}_{old}$ is infeasible (line 2), or both are feasible but ${s}_{new}$ results in a smaller travel distance (line 4), ${s}_{new}$ is accepted. If both are infeasible, the algorithm compares the violation scores of both solutions (line 8) and selects the one with the lower score. This approach allows infeasible solutions to evolve gradually toward feasibility while minimizing the overall travel distance.
Algorithm 1 Constraint-Aware Heuristic
@@ 3.2.4 Constraint-Aware Heuristic
1:  if Checker(${s}_{new}$) is feasible then
2:     if Checker(${s}_{old}$) is infeasible then
@@
4:     else if Cost(${s}_{new}$) $<$ Cost(${s}_{old}$then
5:        return true;
6:     end if
7:  else if Checker(${s}_{old}$) is infeasible then
8:     if Scorer(${s}_{new}$) $<$ Scorer(${s}_{old}$then
9:        return true;
10:     end if
11:  end if
@@ 3.3 Augmented Heuristic Solver
The solver has a general single-point-based search backbone heuristic framework. The automatically generated constraint handling codes are used for this backbone heuristic framework.
3.3.1 Backbone Heuristic Framework
The backbone heuristic solver mainly consists of 1) destroy&repair and 2) local search.
In the destroy phase, we employ multiple operators, including random removal and string removal (Christiaens & Vanden Berghe, 2020), to selectively remove customers or partial routes from the current solution. The choice of destroy operators is determined using a roulette wheel selection mechanism, which assigns higher probabilities to operators that performed well in previous iterations. The repair phase reinserts removed customers into the solution using a greedy repair operator, aiming to construct a feasible solution with shorter routes. Specific details of these operators are provided in the Appendix B.
Following the destroy and repair phases, the solution undergoes a local search process to further refine its quality. We utilize a set of local search operators, including 2-OPT, SWAP, and SHIFT, and the best solution found among these operators is selected. To avoid premature convergence, the Record-to-Record Travel (RRT) criterion is applied (Dueck, 1993; Santini et al., 2018), allowing the acceptance of slightly worse solutions within a predefined threshold, thus maintaining a balance between intensification and diversification.
The entire process is iteratively repeated until a termination condition is met, such as a time limit.
3.3.2 Constraint-Aware Heuristic
@@
Table 1: A classification of common VRP variants is presented based on six constraint types, with abbreviations provided for both the constraints and VRP variants in parentheses. Each constraint type includes four distinct variants, with examples provided for each, resulting in a total of 24 constraints. All of these constraints have already attracted attention from researchers.
Basic Constraints VRP Variants
Vehicle Capacity (C) Capacitated VRP (CVRP) (Vidal, 2022; Luo et al., 2023)
Heterogeneous CVRP (HCVRP) (Lai et al., 2016)
Multi-Product VRP (MVRP) (Yuceer, 1997)
Dynamic CVRP (DCVRP) (Powell, 1986)
Distance Limit (L) VRP with Distance Limit (VRPL) (Laporte et al., 1985; Li et al., 1992)
Heterogeneous VRPL (HVRPL) (Lee et al., 2021)
Recharging VRP (RVRP) (Conrad & Figliozzi, 2011; Erdoğan & Miller-Hooks, 2012)
Dynamic VRPL (DVRPL) (Suzuki, 2011; Khouadjia et al., 2012; Qian & Eglese, 2016)
Time Windows (TW) VRP with time windows (VRPTW) (Solomon, 1987)
Heterogeneous VRPTW (HVRPTW) (Ren et al., 2010; Vidal et al., 2014)
VRP with Multiple Time Windows (VRPMTW) (Belhaiza et al., 2014)
Dynamic VRPTW (DVRPTW) (Ghiani et al., 2003; Pillac et al., 2013)
Pickup and Delivery (PD) VRP with mixed pickup and delivery (VRPMPD) (Avci & Topaloglu, 2015)
Heterogeneous VRPMPD (HVRPMPD) (Avci & Topaloglu, 2016)
Multi-Product VRPMPD (MVRPMPD) (Zhang et al., 2019)
Dynamic VRPMPD (DVRPMPD) (Gendreau et al., 2006)
Same Vehicle (S) VRP with Same Vehicle Constraint (VRPSVC) (Kumar & Panneerselvam, 2012)
Clustered VRP (CluVRP) (Battarra et al., 2014; Islam et al., 2021)
VRP with Sequential Ordering (VRPSO) (Escudero, 1988; Gambardella & Dorigo, 2000)
VRP with Incompatible Loading Constraint (VRPILC) (Wang et al., 2015)
Priority (P) Precedence constrained VRP (PVRP) (Kubo & Kasugai, 1991)
VRP with Relaxed Priority Rules (VRPRP) (Doan et al., 2021)
VRP with Multiple Priorities (VRPMP) (Yang et al., 2015)
VRP with d-Relaxed Priority Rule (VRP-dRP) (Dasari & Singh, 2023; dos Santos Teixeira & de Araujo, 2024)
@@ 4 RoutBench
Recent advancements in LLMs have opened up new possibilities for automatically generating routing solvers to address different VRP variants (Xiao et al., 2023). These methods have shown promising automation capabilities and have been applied to a few VRP variants (Chen et al., 2023). However, their evaluations are often constrained to some simple problems, leaving a significant gap in assessing their generalization ability. There is yet to be a VRP benchmark that can evaluate the generalization ability of these LLM-based methods, especially under complex yet common constraints in real-world scenarios.
Thus, we propose RoutBench, a benchmark dataset that includes 1,000 VRP variants derived from 24 constraints, each has three different sizes: 25, 50, and 100. As shown in Table 1, 4 variants are selected for each of the six constraint types: Vehicle Capacity (C), Distance Limits (L), Time Windows (TW), Pickup and Delivery (PD), Same Vehicle (S) and Priority (P). These constraint types are representative due to their practical significance and theoretical challenges, which already have attracted researchers’ attention. These 1,000 problem instances can be divided into simple and hard based on the number of constraints.
  • RoutBench-S: 500 simple VRP variants with 3 or fewer constraints
  • RoutBench-H: 500 hard VRP variants with 4 or more constraints
Each problem instance is comprised of three components: 1) the problem description, which is a natural language explanation of the problem; 2) the instance data, including the geometric positions of nodes and the constraint parameters; and 3) the validation code, used to confirm whether a solution adheres to user requirements and satisfies all constraints.
By offering a comprehensive and standard dataset for VRP variants with diverse and practical constraints, RoutBench provides a valuable test benchmark for advancing research on vehicle routing algorithms, fostering the development of practical routing solvers.
@@ 4.1 Dataset Construction
The RoutBench is constructed by combinations of six basic constraint types: vehicle capacity, distance limits, time windows, pickup and delivery, same vehicle constraints, and node priority. We design four representative variants derived from each basic constraint type that incorporate variations such as heterogeneous vehicle fleets, multidimensional resource limits, dynamic changes, and others. To produce one problem instance, we first pick one basic constraint combination and then pick one real constrain variant for each chosen basic constraint type. Notice that Vehicle Capacity (C) and Pickup and Delivery (PD) are mutually exclusive, since PD is essentially a derivative of C. The total number of problem combinations is :
@@
where $\binom{6}{k}$ represents the number of ways to select $k$ constraints from the six types, and $4^{k}$ accounts for the four variations per constraint. After excluding combinations where vehicle capacity and pickup and delivery coexist, the total number of feasible combinations is reduced to 5624. From these, 1000 unique problem instances are uniformly sampled based on the order of all combinations, ensuring even coverage across the solution space. To balance complexity, 500 problems are sampled with three or fewer constraints, while the remaining 500 include more than three constraints.
Each problem instance includes detailed parameters such as customer locations, demands, time windows, and other relevant attributes, with data generated using the Solomon C103 dataset as a base. Problem sizes include 25, 50, and 100 nodes. To ensure every problem has at least one feasible solution, the VRP solver in this study directly incorporates a demand verification code into its Selection Strategy. This strategy guides the search process during problem generation, ensuring that each problem is constructed with a valid and feasible solution.
The detailed descriptions of the 24 problem instances are provided in Appendix D. These examples illustrate the diversity of problem settings and serve as a representative subset of the dataset’s broader scope. By systematically combining constraints, leveraging validation mechanisms, and ensuring feasibility, RoutBench offers a diverse and reliable dataset for benchmarking VRP solvers.

### v3

@@ Abstract
Real-world Vehicle Routing Problems (VRPs) are characterized by a variety of practical constraints, making manual solver design both knowledge-intensive and time-consuming. Although there is increasing effort in automating the design of routing solvers, existing research has explored only a limited array of VRP variants and fails to adequately address the complex and prevalent constraints encountered in real-world situations. We propose Automatic Routing Solver (ARS), which leverages Large Language Model (LLM) agents to enhance a backbone metaheuristic framework. ARS automatically generates constraint-aware heuristic code using problem descriptions and representative constraints selected from a database, offering greater flexibility and interpretability than existing methods without relying on cumbersome modeling rules. Alongside ARS, we introduce RoutBench  [^1] https://github.com/Ahalikai/ARS-Routbench/, a benchmark comprising 1,000 VRP variants derived from 24 attributes, designed to rigorously evaluate the effectiveness of automatic routing solvers in handling VRPs with diverse practical constraints. We evaluate ARS against seven LLM-based methods on both common VRPs and RoutBench. Experimental results show that ARS achieves a success rate of over 90% on common VRPs and over 60% on RoutBench, outperforming other methods by at least 30% in success rate. In terms of solving performance, ARS also significantly surpasses general-purpose solvers, demonstrating higher efficiency across many VRP variants.
@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a fundamental Combinatorial Optimization Problem (COP) that plays a critical role in logistics, transportation, manufacturing, retail distribution, and delivery planning [1, 2]. In these scenarios, the objective of VRPs is to efficiently allocate and plan vehicle routes to meet various requirements while minimizing overall routing costs. These requirements often include constraints such as vehicle capacities, time windows, and duration limits, resulting in numerous variants of VRPs in practical applications [3]. However, existing heuristics are usually problem-specific. When the problem changes slightly (e.g., a minor modification to the requirements), a lot of effort is required for experts to redesign the heuristic to make it effective in solving the new problems [4, 5, 6].
Large Language Models (LLMs) have shown powerful reasoning and code-generation capabilities [7, 8, 9]. By integrating these functionalities, users can express their specific requirements in natural language, enabling the models to automatically design algorithms to address VRPs [10]. Most existing works primarily leverage the LLM to solve a small number of standard VRPs [11, 12] and cannot be applied to complex VRPs. Some recent studies have employed LLMs to formulate routing problems as integer programming models and solve them using general-purpose solvers [13, 14]. However, these approaches may face challenges in adhering to standard modeling practices, and the solvers often exhibit low efficiency, limiting their ability to handle complex real-world constraints.
@@
  • We comprehensively validate our approach on commonly-used problems and RoutBench. The results show that ARS can automatically handle 91.67% of common problems, achieving a significant improvement across all benchmarks compared to other LLM-based methods and general solvers.
@@ 2 Problem Formulation
where $\mathbf{x}=\{x_{ij}\mid i,j\in\mathcal{V},\ i\neq j\}$ represents the set of decision variables, $x_{ij}$ is a binary variable that indicates whether the route from $i$ to $j$ is used. The feasible solution space $C$ is defined by constraints. In this paper, we consider VRPs with a variety of real-world constraints such as vehicle capacity, travel distance, and time windows, thereby extending the basic VRP, as seen in the Capacitated VRP (CVRP) [1] and the VRP with Time Windows (VRPTW) [15]. Moreover, new constraints often emerge in real-world scenarios. For example, VRP variants related to vehicle capacity include Heterogeneous VRP (HVRP), which considers vehicles with different capacities [16], Multi-Product VRP (MPVRP), which addresses the need to transport multiple types of products [17], and dynamic demands [18]. VRP variants that incorporate these real-world constraints are more prevalent and practically significant in real-world applications.
@@ 3.1 Database
  • $C_{k}$: The constraint feasibility checking program. It checks whether a solution belongs to the feasible solution space described by $I_{k}$. The program is given a solution and returns "True" if the corresponding constraint is satisfied.
@@ 3.2 Constraint-Aware Heuristic
We first select relative constraints from the database. Then we sequentially generate the constraint checking and scoring programs for the target VRP variants, given the selected constraints and the problem description. Finally, the constraint-aware heuristic is generated based on the designed constraint checking and scoring programs.
@@ 3.2.1 Constraint Selection
Given an input problem description $I$, we instruct LLM agents to automatically select a subset of relevant constraints $S$ from the database $D$ to provide reference information for subsequent constraint program generation. There can be two cases. In the first case, LLM agents select one or more relative constraint examples. In the second case, if no constraints are recognized as related to the input $I$, the base case $(I_{0},C_{0})$, with no additional constraints, is selected.
This step is similar to Retrieval-Augmented Generation (RAG) methods that retrieve information from external sources [19, 20], while our candidate reference information is a pre-defined database with representative constraints.
@@ 3.2.3 Constraint Scoring Program Generation
In practice, heuristic algorithms operate within a variable space that is typically much larger than the feasible solution space. This discrepancy presents significant challenges in identifying feasible solutions. To address this, heuristic methods may permit the presence of infeasible solutions during the search process, as this allows for the evaluation of potential improvements in solution quality [21, 22].
@@
3.2.4 Constraint-Aware Heuristic Generation
We present the Constraint-Aware Heuristic (CAH) based on the constraint checking (Checker) and scoring programs (Scorer). As seen in Algorithm 1, the constraint handling method evaluates whether a new solution ${s}_{new}$ improves upon an old solution ${s}_{old}$. It first verifies whether ${s}_{old}$ is feasible (line 1). If ${s}_{old}$ is feasible, it then verifies whether ${s}_{new}$ is feasible and has a smaller travel distance (line 2). If both conditions are satisfied, ${s}_{new}$ is accepted. If ${s}_{old}$ is infeasible, but ${s}_{new}$ is either feasible or has a lower violation score than ${s}_{old}$ (line 5), ${s}_{new}$ is also accepted. This approach allows infeasible solutions to evolve gradually toward feasibility while minimizing the overall travel distance.
Algorithm 1 Constraint-Aware Heuristic (CAH)
@@ 3.2.4 Constraint-Aware Heuristic Generation
1:  if Checker(${s}_{old}$) is feasible then
2:     if Checker(${s}_{new}$) is feasible and Cost(${s}_{new}$) $<$ Cost(${s}_{old}$then
@@
4:     end if
5:  else if Checker(${s}_{new}$) is feasible or Scorer(${s}_{new}$) $<$ Scorer(${s}_{old}$then
6:     return true;
7:  end if
8:  return false;
@@ 3.3 Augmented Heuristic Solver
The solver has a general single-point-based search backbone heuristic framework, which utilizes automatically generated constraint-aware heuristics to solve various VRP variants. This backbone heuristic solver mainly consists of 1) destroy&repair and 2) local search.
In the destroy phase, we employ multiple operators, including random removal and string removal [23], to selectively remove customers or partial routes from the current solution. The choice of destroy operators is determined using a roulette wheel selection mechanism, which assigns higher probabilities to operators that performed well in previous iterations. The repair phase reinserts removed customers into the solution using a greedy repair operator, aiming to construct a feasible solution with shorter routes. Specific details of these operators are provided in Appendix B.3.
Following the destroy and repair phases, the solution undergoes a local search process to further refine its quality. We utilize a set of local search operators, including 2-OPT [24], SWAP [25], and SHIFT [26], and the best solution found among these operators is selected. To avoid premature convergence, the Record-to-Record Travel (RRT) criterion is applied [27, 28], allowing the acceptance of slightly worse solutions within a predefined threshold, thus maintaining a balance between intensification and diversification. The entire process is iteratively repeated until a termination condition is met, such as a time limit.
@@ 4 RoutBench
Recent advancements in LLMs have opened up new possibilities for automatically generating routing solvers to address different VRP variants [13, 29]. However, they are typically evaluated on only a few dozen simple problems, leaving a significant gap in assessing their generalization ability. There is yet to be a VRP benchmark that can evaluate the generalization ability of these LLM-based methods, especially under complex yet common constraints in real-world scenarios.
Table 1: A classification of common VRP variants is presented based on six constraint types, with abbreviations provided for both the constraints and VRP variants in parentheses. Each constraint type includes four distinct variants, with examples provided for each, resulting in a total of 24 constraints.
Basic Constraints VRP Variants Basic Constraints VRP Variants
Vehicle Capacity (C) Capacitated VRP (CVRP) [30, 31] Distance Limit (L) VRP with Distance Limit (VRPL) [32, 33]
Heterogeneous CVRP (HCVRP) [16] Heterogeneous VRPL (HVRPL) [34]
Multi-Product VRP (MVRP) [17] Recharging VRP (RVRP) [35, 36]
Dynamic CVRP (DCVRP) [18] Dynamic VRPL (DVRPL) [37, 38, 39]
Time Windows (TW) VRP with Time Windows (VRPTW) [15] Pickup and Delivery (PD) VRP with Mixed Pickup and Delivery (VRPMPD) [40]
Heterogeneous VRPTW (HVRPTW) [41, 42] Heterogeneous VRPMPD (HVRPMPD) [43]
VRP with Multiple Time Windows (VRPMTW) [44] Multi-Product VRPMPD (MVRPMPD) [45]
Dynamic VRPTW (DVRPTW) [46, 47] Dynamic VRPMPD (DVRPMPD) [48]
Same Vehicle (S) VRP with Same Vehicle Constraint (VRPSVC) [49] Priority (P) Precedence constrained VRP (PVRP) [50]
Clustered VRP (CluVRP) [51, 52] VRP with Relaxed Priority Rules (VRPRP) [53]
VRP with Sequential Ordering (VRPSO) [54, 55] VRP with Multiple Priorities (VRPMP) [56]
VRP with Incompatible Loading Constraint (VRPILC) [57] VRP with d-Relaxed Priority Rule (VRP-dRP) [58, 59]
Thus, we propose RoutBench, a benchmark dataset that includes 1,000 VRP variants derived from 24 constraints. These constraints are chosen for their practical significance and theoretical challenges, as highlighted in Table 1. This design not only expands the test scale by two orders of magnitude but also provides an opportunity to evaluate algorithms on unseen VRPs. If an algorithm can effectively solve these unseen problems, it demonstrates the potential to address new real-world challenges, better meeting practical application needs.
Figure 2: Analysis of problem distribution across common problems, RoutBench, and all problems.
Figure 3: The frequency of constraint usage is analyzed for RoutBench-S and RoutBench-H.
@@ 4.1 Dataset Construction
The RoutBench is constructed by combinations of six basic constraint types: Vehicle Capacity (C), Distance Limits (L), Time Windows (TW), Pickup and Delivery (PD), Same Vehicle (S), and Priority (P). We design four representative variants derived from each basic constraint type that incorporate variations such as heterogeneous vehicle fleets, multidimensional resource limits, dynamic changes, and others. To produce one problem instance, we first pick one basic constraint combination and then pick one real constraint variant for each chosen basic constraint type. Notice that Vehicle Capacity (C) and Pickup and Delivery (PD) are mutually exclusive, since PD is essentially a derivative of C. The total number of problem combinations is :
@@
where $\binom{6}{k}$ represents the number of ways to select $k$ constraints from the six types, and $4^{k}$ accounts for the four variations per constraint. After excluding combinations where vehicle capacity and pickup and delivery coexist, the total number of feasible combinations is reduced to 5624. From these, 1000 unique problem instances are uniformly sampled based on the order of all combinations, ensuring even coverage across the solution space. To balance complexity, RoutBench-S includes 500 problems with three or fewer constraints, while RoutBench-H consists of the remaining 500 problems with more than three constraints.
Each problem instance is comprised of three components: 1) the problem description, which is a natural language explanation of the problem; 2) the instance data, including the geometric positions of nodes and the constraint parameters, with data generated using the Solomon C103 dataset [15] as a base; and 3) the validation code, used to confirm whether a solution adheres to user requirements and satisfies all constraints. Problem sizes include 25, 50, and 100 nodes. All instances in RoutBench come with Best Known Solutions (BKS), with details provided in Appendix D.5.
The detailed descriptions of the 24 problem instances are provided in Table 6. These examples illustrate the diversity of problem settings and serve as a representative subset of the dataset’s broader scope. By systematically combining constraints, leveraging validation mechanisms, and ensuring feasibility, RoutBench offers a diverse and reliable dataset for benchmarking VRP solvers.